OpenStudy (anonymous):

How do I evaluate: 5^-4 - 5^-6/5^-3 + 5^-5?

jhonyy9 (jhonyy9):

\[1/5^{4} -(1/5^{6})/1/5^{3} +1/5^{5}=1/5^{4} -(1/5^{6-3}) +1/5^{5}\]

jhonyy9 (jhonyy9):

\[1/5^{4} +1/5^{5} -1/5^{3}=(5+1)/5^{5} -1/5^{3}\]

jhonyy9 (jhonyy9):

\[6/5^{5} -1/5^{3} = (6-5^{2})/5^{5} = -19/5^{5}\]